OpenStudy (anonymous):

PLEASE HELP --> WILL GIVE MEDAL --> PLEASE EXPLAIN Task 3 Part 1: Create a scenario for an arithmetic sequence. For example: Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. __________________________________________________________ Part 2: Create a scenario for a geometric sequence. For example: Anthony goes to the gym for ______ minutes on Monday. Every day he _________his gym time by ____________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? __________________________________________________________ Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

Jasmine practices the piano for _10_____ minutes on Monday. Every day she ____increases_______ her practice time by __15 minutes_______. If she continues this pattern, how many minutes will she practice on the 7th day? Use your scenario to write the function for the 7th term in your sequence using sequence notation. Monday- 10 minutes Tuesday - 25 minutes Wednesday - 40 minutes Thursday - 55 minutes Friday - 70 minutes Saturday - 85 minutes Sunday - 100 minutes She will practice piano for 100 minutes on the 7th day, Sunday. Function: An = a1 + d(n – 1) Day N = 10 +15(n-1) Day 7= 10+15 (7-1) Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10%________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. (I rounded the decimals here) Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes He will spend 29.282 minutes at the gym on the 5th day, Friday. Sn=a*1−r^n/1−r where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. S(5)=20 1-1.1^5 / 1-1.1 =122.102